OpenStudy (deadshot):

What is the slope of the line that is perpendicular to y = 2x – 3?

hartnn (hartnn):

can you find slope of this line ? y= 2x-3 ?

OpenStudy (deadshot):

the slope is 0 for y=2x-3

hartnn (hartnn):

nopes, how ? compare it with, y= mx+c where m= slope

OpenStudy (deadshot):

oh, so then the slope is 2?

hartnn (hartnn):

yes now product of slope of perpendicular lines =-1 so, to get the slope of perpendicular line, you just take negative reciprocal of this slope

OpenStudy (deadshot):

so then the slope of the perpendicular line would be -2?

OpenStudy (anonymous):

\[ 0=\begin{bmatrix} 1 \\ m \end{bmatrix} \cdot \begin{bmatrix} 1 \\ m' \end{bmatrix} \implies 1 + mm' = 0 \implies mm' = -1 \implies m' = -\frac 1 m \]

hartnn (hartnn):

you forgot to do reciprocal ?

OpenStudy (deadshot):

yeah, It's been a while since I've done this

OpenStudy (anonymous):

So if \(m\) is the slope of a line, a line who is perpendicular has slope \(m' = -\frac 1 m\).

hartnn (hartnn):

what is the reciprocal of 2 ?

OpenStudy (deadshot):

ok, so then the reciprocal of 2 would be \[- \frac{ 1 }{ 2 }\] right?

hartnn (hartnn):

reciprocal of 2 is 1/2 so, the slope of line will be -1/2 yes, correct.

hartnn (hartnn):

slope of perpendicular line**

OpenStudy (deadshot):

Okay, thanks!

hartnn (hartnn):

welcome ^_^